solve the system by substitution.
{-x+y=-1
{3x-2y=1

Answer by deborabr(181)

{-x+y=-1
{3x-2y=1
.
{-x+y=-1
-x=-1-y
x=1+y
.
{3x-2y=1
3(1+y)-2y=1
3+3y-2y=1
3+y=1
y=1-3
y=-2
.
{-x+y=-1
-x+(-2)=-1
-x-2=-1
-x=-1+2
-x=1 *(-1)
x=-1
.
y=-2 and x=-1